Two slits are separated by 0.320 mm. A beam of 500nm light strikes the slits, producing an interference pattern. Determine the number of maxima observed in the angular range -30.0degree < s < 30.0degree, where s is an angle


1
Expert's answer
2021-02-22T16:04:52-0500


The slit separation is d = 0.320 mm = 0.320"\times" 10-3 m


The wavelength of the beam is λ = 500 nm = 500*10-9 m


Let there be n fringes (bright& dark) of width "x" on one side of the central bright.


Then the total distance on the screen within the angular separation is Y = "nx"


Then,

"tan\theta= \dfrac{Y}{D}=\dfrac{nx}{D}"


"nx= Dtan\theta"

Where D is the separation of slits and the screen.

The width of each fringe is

"\Rightarrow x=\dfrac{\lambda D}{d}"


"\Rightarrow nx=\dfrac{n\lambda D}{d}"


"\Rightarrow Dtan\theta=\dfrac{n\lambda D}{d}"


"\Rightarrow n= \dfrac{d\times tan\theta}{\lambda}"

"=\dfrac{(0.320\times 10^{-3})(tan30)}{500\times 10^{-9}}"


"=369.5"

The total number fringes in the angular separation of "-30^\degree"to "30^\degree"is 369.5"\times"2 = 739

The total number of bright fringes(including the central bright) is 370